# Scalable Oversight for Superhuman AI via Recursive Self-Critiquing

### v3

@@ 3.2 Setup
$C^{3}$ 67.19 67.19 60.16 3.766 14.23
GAOKAO Math Response 66.29 85.94 - 3.194 14.58
@@
$C^{3}$ 93.75 93.75 87.50 4.031 14.14
Execution
We write standardized guidelines for all tasks through descriptive instructions and examples, detailed in Appendix A. Tasks are distributed in data packages with specified submission deadlines, with random assignments across different critique levels to ensure all annotators participate in all stages. To maintain efficiency, we set a 20-minute time limit for each question at every stage. While the time constraint exists, it is managed through package-level deadlines to allow flexible time allocation. Annotators complete a set number of tasks daily within their scheduled working hours. Feedback sessions are held to collect comments and suggestions for procedure or guideline improvement. We assign personnel for process management and annotation quality assurance.
Metrics
We assess the effectiveness of recursive criticism through three metrics: (i) accuracy measures the consistency with the standard answers; (ii) completion time records the duration of the entire process, including analysis and answer provision; (iii) confidence reflects self-assessed certainty in the final answer on a five-point scale.
3.3 Critique of Critique can be Easier than Critique
We validate the hypothesis that critique of critique is easier than critique across five tasks. The results in Table 1 present consistent improvements from response to critique to $C^{2}$ stages. Take GAOKAO Math as an example, the average accuracy improves from 66.29% (response) to 82.50% (critique) and further to 90.62% ($C^{2}$), while completion time remains stable or slightly decreases (e.g., from 18.36 to 15.82 minutes for CET 6). Under comparable effort, majority voting shows similar trends. For example, accuracy improves from 81.81% (response) through 86.61% (critique) to 90.62% ($C^{2}$) in GAOKAO Math, demonstrating the advantage of higher-order criticism. Compared to naive voting, average accuracy consistently performs better. Take GAOKAO Math as an example, naive voting achieves only 66.41% at critique stage and 81.25% at $C^{2}$, significantly lower than average accuracy 90.62%. These results validate that recursive critique generates new insights rather than merely aggregating previous judgments. Moreover, annotator confidence shows steady improvement across stages, suggesting that higher-order criticism becomes more tractable.
3.4 Recursive Critique Remains Consistently Easier
We extend the recursive criticism to the third-order critique ($C^{3}$) on two representative tasks. As shown in Table 2, accuracy improves continuously at the $C^{3}$ level in both tasks, with CET-6 increasing from 60.94% at $C^{2}$ to 67.19%, and GAOKAO Math from 90.62% to 93.75%. More importantly, under comparable computational effort, majority voting shows similar improvements, reaching 67.19% for CET-6 and 93.75% for GAOKAO Math at $C^{3}$ level. Furthermore, naive voting achieves substantially lower performance than regular accuracy. Meanwhile, confidence scores improve as completion time decreases. These results demonstrate that recursive critique remains consistently easier and extends beyond mere computational scaling or consensus aggregation.
@@ 4 Can Recursive Self-Critiquing Enable Human Oversight of AI?
Table 3: Performance comparison across recursive criticism stages, with human accuracy subscripts showing difference from previous-stage AI accuracy. Results from Qwen2.5-7B/72B-Instruct on mathematics and English tests, including accuracy, confidence (1-5), and completion time (minutes).
Dataset Stage Human Accuracy AI Accuracy Confidence Time(min)
@@
$C^{3}$ 71.88+3.91 - 3.813 16.86
4.1 Tasks
We select tasks where AI demonstrates superior performance compared to our human annotators’ capabilities. Specifically, the GAOKAO Math comprises the last two multiple-choice questions from the high school mathematics examination (Zhang et al., 2023), while the TEM4 (Test for English Majors Grade Four) includes reading comprehension questions. Importantly, these mathematics questions demand advanced problem-solving skills, and similarly, TEM-4 questions require professional-level English proficiency. As a result, both tasks are beyond most annotators’ abilities.
4.2 Setup
We employ the same annotators and evaluation metrics as in Human-Human experiments. To generate AI responses, we utilize both Qwen-7B-Instruct and Qwen-72B-Instruct models (Qwen et al., 2025) to examine recursive criticism across different AI capability levels. We filter out questions on which models achieve either 0% or 100% accuracy, as these extremely easy or difficult tasks lead to uniform model outputs, making them unsuitable for validating recursive criticism. For each question, the AI model first outputs initial responses, then performs self-criticism recursively to certain orders ($C^{1}$, $C^{2}$). Human annotators only evaluate the final AI outputs. The annotation process follows the Human-Human procedure, except outputs are AI-generated.
4.3 Experimental Results
@@
Table 3 indicates that humans’ response accuracy is lower than that of AI, showing that AI surpasses humans in these tasks. However, for the following critique, humans consistently achieve higher accuracy than AI’s previous outputs. For example, with Qwen2.5-7B on GAOKAO Math, human accuracy reaches 53.12% at the critique stage (7.03% higher than AI’s initial 46.09%), and further increases to 56.25% at $C^{2}$ (8.59% above AI’s critique stage). This finding suggests that recursive criticism enables human supervision even when AI outperforms humans for generation.
@@
Despite processing more information at higher levels, completion time generally decreases or remains stable. For TEM4 with Qwen-72B, time decreases from 22.44 minutes at response level to 15.47 minutes at $C^{2}$. Meanwhile, annotator confidence shows consistent improvement across levels and model scales, particularly in the mathematics task with Qwen-72B where confidence increases from 2.19 to 3.63. These results suggest that recursive criticism makes evaluation more tractable.
@@ 5.1 Setup
Model Preperation
We investigate the dynamics of supervisory effectiveness across varied pairings of supervised and critic models with different capability levels. To establish our experimental framework, we utilize the Qwen2.5 series models (Qwen et al., 2025), operating under the established premise that model capability generally correlates with parameter size. However, different variants of the Qwen2.5-instruct series may have undergone different post-training procedures, we implemented a standardization approach. Specifically, we randomly sampled 282k instances from the open-source TULU-3-SFT dataset (Lambert et al., 2024) and fine-tuned the Qwen2.5-base model series.
@@
For the unambiguous measurement of supervision quality, we select mathematical tasks due to their objective verifiability. The experimental data were drawn from the DeepScaleR dataset (Luo et al., 2025), with 512 randomly sampled instances constituting the test set and the remainder serving as training data. Throughout our experimental validation, we employ the math-verify library (Kydlíček and Gandenberger, 2025) to obtain reliable ground truth signals.
@@
In our experiments, the supervised model initially engages in recursive self-critiquing at varying orders. Subsequently, the critic model conducts a final higher-order critique based on the supervised model’s outputs. Details of prompts and sampling strategies are provided in the Appendix B. Adhering to established methodologies in RLHF (Ouyang et al., 2022), we leverage these final critiques to construct preference data and train reward models. To mitigate potential confounding effects stemming from architectural similarities between the reward and SFT models, we select Llama3.1-8B (Meta, 2024) as the foundation for the reward model. The resultant reward model will be used for Best-of-N sampling, enabling systematic evaluation of supervisory efficacy across diverse model-critic combinations.
@@
To quantify supervision effectiveness, we adopt the performance recovered (PR) metric in accordance with the framework established by Burns et al. (2023):
@@
In this formulation, $x\sim\mathcal{D}$ denotes inputs sampled from distribution $\mathcal{D}$, while $\{y_{i}\}_{i=1}^{n}\sim M(\cdot|x)$ represents $n$ samples generated by model $M$ given input $x$. The learned reward function is expressed as $r(x,y)$, with $r^{*}(x,y)$ designating the ground truth reward function. Within the context of mathematical tasks, where $r^{*}$ serves as a binary correctness indicator, this ratio effectively captures the accuracy of reward-guided Best-of-N sampling relative to pass@N.
5.2 Experiment Results
Figures 3 and 3 illustrate our experimental results under two settings: (1) A constant-size critic model evaluating supervised models of varying capacities, where each supervised model performs recursive self-critiquing before receiving the critic’s final assessment. (2) A fixed-size supervised model evaluated by critics of varying capabilities, where the supervised model completes its recursive self-critiquing prior to the critic’s final evaluation.
Figure 2: Performance Recovered scores with a fixed 7B critic model and supervised models of varying sizes.
Figure 3: Performance Recovered scores with a fixed 7B supervised model and critic models of varying sizes.
Recursive self-critiquing benefits in weak-to-strong supervision, while direct supervision exhibits superior performance in strong-to-weak settings.
Figure 3 demonstrates that when supervised models are larger than the 7B critic model, higher-order critiques generally yield improved performance compared to lower-order critiques. Conversely, when critic models are stronger than the supervised model, direct critique produces better results than allowing the supervised model to engage in higher-order critiquing. Figure 3 illustrates similar phenomena. This asymmetry may be explained by the fact that self-critiquing from weaker models is not necessarily effective and may even be misleading for stronger model supervision. In contrast, when supervising stronger models, recursive self-critiquing by stronger models can provide beneficial signals. These findings support recursive self-critiquing as a promising approach to scalable oversight, particularly in scenarios where humans (as the "weaker model") oversee increasingly capable AI systems (the stronger model). We provide additional exploration of AI self-critiquing in Appendix C.
6 Discussion
@@
RLHF has emerged as the dominant approach in AI alignment, building upon the principle that "verification is easier than generation" (Irving et al., 2018b). However, the optimal RLHF setup requires direct human preferences for optimization, which necessitates the deployment of static reward models as proxies due to challenges in acquiring real-time human feedback. Such reliance on static proxies introduces reward hacking (Gao et al., 2022; Karwowski et al., 2023); optimizing against these models rather than ideal human preferences leads to policies that diverge from intended objectives due to Goodhart’s Law (Manheim and Garrabrant, 2019; Karwowski et al., 2023; Wen et al., 2024). While approaches such as iterative annotation and tool augmentation (Li et al., 2024a; Gou et al., 2024) provide intermediate solutions, they ultimately face limitations in supervision capability. The recursive criticism framework offers a promising approach by enabling sustained human oversight even as direct evaluation becomes intractable.
@@
The effectiveness of recursive self-critiquing stems from several key mechanisms. Higher-order criticism progressively shifts attention from specific details to abstract evaluation principles, making complex evaluations more tractable. Each critique level provides structured context for subsequent analyses, while the recursive structure transforms absolute tasks into pairwise judgments, leveraging humans’ cognitive advantage in relative assessment over absolute evaluation (Jones and Inglis, 2015; Kelly et al., 2022). Despite these advantages, our further AI-AI experiments in Appendix C suggest current models may lack sufficient critique capabilities, particularly in identifying critical errors (Xi et al., 2024), likely due to the sparsity of critique data in both pretraining and posttraining. Future work may focus on enhancing model critique capabilities (Wang et al., 2024a; Yu et al., 2025; Ankner et al., 2024).
7 Related Work
Reinforcement Learning from Human Feedback (Ouyang et al., 2022) has emerged as a foundational approach for aligning AI systems with human preferences. However, as AI capabilities exceed human expertise in certain domains, humans may no longer provide effective supervision signals (Amodei et al., 2016). To respond to this limitation, several works explore potential methodologies to enable weak annotators to supervise strong AI systems (Burns et al., 2023). The debate protocol (Irving et al., 2018a) involves agents arguing for opposing answers, with studies showing promising results (Khan et al., 2024; Michael et al., 2023) despite some limitations (Kenton et al., 2024a). Unlike debate’s zero-sum framework, our approach assumes higher-order critic tasks are easier. Task decomposition (Christiano et al., 2018; Wu et al., 2021) breaks complex oversight into manageable sub-problems, though our method employs depth-first rather than breadth-first search in problem decomposition. Our majority vote baseline builds on self-consistency methods (Wang et al., 2023), which enables superhuman model evaluation through consistency checks (Fluri et al., 2023).
@@ 8 Conclusion
This work investigates how to obtain reliable supervision signals when AI capabilities surpass human abilities. Through comprehensive experiments in Human-Human, Human-AI, and AI-AI contexts, we examine the hypotheses that critique of critique is easier than critique and demonstrate that this difficulty relation holds recursively. The experiments demonstrate the potential of recursive self-critiquing mechanisms for maintaining effective oversight in scenarios where direct human evaluation becomes infeasible, and suggest a promising pathway for scalable oversight.
@@ Appendix A Human Experiments Guidelines
This section details the guidelines and quality assurance of involved in the Human-Human and Human-AI experiments. We establish consistent and comprehensive guidelines for annotation tasks at different stages across different tasks.
Our guidelines emphasize the quality of reasoning process over accuracy rates, requiring annotators to clearly articulate their thinking process without accessing external references. While accuracy is encouraged, the primary focus is on providing clear, well-reasoned justifications for their decisions. Annotators are instructed to invest their time primarily in analytical thinking, expressing their reasoning in clear, concise, and logically coherent natural language. The guidelines provide suggested formats but maintain flexibility, prioritizing the clear documentation of thought processes over rigid adherence to specific forms [^2] Fixed templates were initially tested but abandoned as annotators reported them inflexible and including unnecessary burden.. We provide detailed instruction at each stage in following sections.
@@ A.1 Response Stage
{CJK*}
UTF8gkai
Table 4: High quality and low quality response examples.
Quality Definition Type Example and Translation
Math Origin: 首先化简$f(x)=2\cos^{2}x-\sin^{2}x+2$,根据二倍角公式$\cos 2x=2\cos^{2}x-1$,得到$12\cos^{2}x=\cos 2x+$ $\sin^{2}x+\cos^{2}x=1$,所以$\sin^{2}x=(1-\cos 2x)/2$。最终得到$f(x)=\frac{3}{2}\cos 2x+\frac{5}{2}$。通过周期计算和最值分析,得到答案B。 Translated: First simplify $f(x)=2\cos^{2}x-\sin^{2}x+2$. Using double angle formula $\cos 2x=2\cos^{2}x-1$, we get $12\cos^{2}x=\cos 2x+$ 1. Since $\sin^{2}x+\cos^{2}x=1$, we derive $\sin^{2}x=(1-\cos 2x)/2$. Finally $f(x)=\frac{3}{2}\cos 2x+\frac{5}{2}$. Through period calculation and maximum analysis, we arrive at answer B.
[…]
@@ A.2 Critique Stage Annotation
{CJK*}
UTF8gkai
Table 5: High quality and low quality critique annotation examples.
Question Response1 Response2
High Quality Examples
Low Quality Examples
认可response 1的答案,不同意Response2的答案,答案应该是(B)。
Question Response1 Response2
High Quality Examples
Low Quality Examples
@@ A.3 Higher-Order Critique Stage
Examples of high-quality and low-quality higher-order critiques are presented in Tables 7 and 8.
{CJK*}
UTF8gkai
Question Response 1 Response 2
Critic 1 Critic 2
High Quality Examples
Low Quality Examples
Critc 1/2的答案是对,应该是(B)。
Question Response 1 Response 2
Critic 1 Critic 2
High Quality Examples
Low Quality Examples
Critics 1/2 are correct, the answer should be (B).

### v4

@@ 3.4 Recursive Critique Remains Consistently Easier
$C^{3}$ 67.19 60.16 3.766 14.23
GAOKAO Math Response 66.29 85.94 3.194 14.58
@@
$C^{3}$ 93.75 87.50 4.031 14.14
@@ 4 Can Recursive Self-Critiquing Enable Human Oversight of AI?
4.1 Tasks
We select tasks based on the criterion that humans find them challenging while AI demonstrates reasonable but not perfect performance, creating suitable conditions for meaningful evaluation of human oversight when AI capabilities exceed human performance. Following this criterion, we select two challenging task types for our experiments:
  • GAOKAO Math comprises the last two multiple-choice questions from the high school mathematics examination (Zhang et al., 2023), which demand advanced problem-solving skills and mathematical reasoning abilities.
  • TEM4 (Test for English Majors Grade Four) includes reading comprehension questions that require professional-level English proficiency and complex text analysis capabilities.
Both tasks are beyond most annotators’ abilities while remaining moderately challenging for AI models. We filter out questions where models achieve either 0% or 100% accuracy, as these overly easy or difficult cases produce uniform outputs that are unsuitable for our validation.
4.2 Setup
We employ the same annotators, annotation procedures, and evaluation metrics as in Human-Human experiments. The annotation process follows the Human-Human procedure, with AI outputs replacing human ones. To obtain AI responses, we utilize both Qwen-7B-Instruct and Qwen-72B-Instruct models (Qwen et al., 2025) to examine recursive critique across different AI capability levels. For each question, the AI model first generates initial responses, then performs self-critique recursively at multiple orders ($C^{1}$, $C^{2}$). Human annotators evaluate AI outputs at each corresponding stage, except for the Response stage where humans complete tasks independently without relying on AI outputs.
4.3 Experimental Results
Table 3: Performance comparison across recursive critique stages, with human accuracy subscripts showing difference from previous-stage AI accuracy. Results from Qwen2.5-7B/72B-Instruct on mathematics and English tests, including accuracy, confidence (1-5), and completion time (minutes).
Dataset Stage Human Accuracy AI Accuracy Confidence (1-5) Time (min)
@@ 4.3 Experimental Results
$C^{3}$ 71.88+3.91 3.813 16.86
@@
Table 3 indicates that human response accuracy is lower than AI accuracy, showing that AI surpasses humans in these tasks. However, in subsequent critique stages, humans consistently achieve higher accuracy than AI’s previous outputs. For example, with Qwen2.5-7B on GAOKAO Math, human accuracy reaches 53.12% at the critique stage (7.03% higher than AI’s initial 46.09%), and further increases to 56.25% at $C^{2}$ (8.59% above AI’s critique stage). This finding suggests that recursive critique enables human supervision even in the tasks which AI outperforms humans.
@@
Despite processing more information at higher levels, completion time generally decreases or remains stable. For TEM4 with Qwen-72B, time decreases from 22.44 minutes at the response level to 15.47 minutes at $C^{2}$. Meanwhile, annotator confidence shows consistent improvement across levels and model scales, particularly in the mathematics task with Qwen-72B where confidence increases from 2.19 to 3.63. These results suggest that recursive critique makes evaluation more tractable.
@@ 5.1 Setup
Model Preparation
We investigate the dynamics of supervisory effectiveness across pairings of supervised and critic models with different capability levels. We use models from the Qwen2.5 series (Qwen et al., 2025), operating under the premise that model capability generally correlates with parameter size. However, different variants of the Qwen2.5-Instruct series may have undergone different post-training procedures, which could introduce confounding factors. To ensure fair comparison, we randomly sample 282K instances from the open-source TULU-3-SFT dataset (Lambert et al., 2024) and fine-tune all models from the Qwen2.5-Base series using the same data.
@@
To ensure objective measurement of supervision quality, we select mathematical tasks due to their verifiable nature. The experimental data are drawn from the DeepScaleR dataset (Luo et al., 2025), with 512 randomly sampled instances as the test set and the remainder as training data. We employ the Math-Verify library (Kydlíček and Gandenberger, 2025) to determine answer correctness and obtain reliable ground truth signals.
@@
In our experiments, the supervised model first performs recursive self-critique at varying orders. Subsequently, a critic model conducts a final higher-order critique based on the supervised model’s outputs. Prompts and sampling strategies are detailed in Appendix B. Following established RLHF methodologies (Ouyang et al., 2022), we use these final critiques to construct preference pairs and train reward models. To avoid potential confounding effects from architectural similarities, we select Llama3.1-8B (Meta, 2024)—a different architecture with comparable model capacity—as the foundation for the reward model. This reward model is then used for Best-of-N sampling to systematically evaluate supervisory efficacy across diverse model-critic combinations.
@@
To quantify supervision effectiveness, we adopt the Performance Recovered (PR) metric following the framework established by (Burns et al., 2023):
@@
In this formulation, $x\sim\mathcal{D}$ denotes inputs sampled from distribution $\mathcal{D}$, and $\{y_{i}\}_{i=1}^{n}\sim M(\cdot|x)$ represents $n$ samples generated by model $M$ given input $x$. The learned reward function is denoted as $r(x,y)$, with $r^{*}(x,y)$ representing the ground truth reward function. Intuitively, the numerator measures the ground truth quality of outputs selected by the learned reward model, while the denominator represents the oracle performance with perfect selection. For mathematical tasks where $r^{*}$ indicates binary correctness, PR thus quantifies how close the learned reward model comes to oracle pass@N performance.
Figure 2: PR scores with a fixed 7B critic model and supervised models of varying sizes.
Figure 3: PR scores with a fixed 7B supervised model and critic models of varying sizes.
5.2 Experimental Results
Figures 3 and 3 present our experimental results under two settings: (1) Figure 3 shows results where supervised models of varying sizes first perform recursive self-critique, followed by evaluation from a fixed 7B critic model at each stage. The critic’s judgments are used to train reward models specific to each supervised model size, which then guide Best-of-N sampling on the corresponding supervised models. The PR metric compares this Best-of-N performance against the oracle Pass@N performance for each model size. (2) Figure 3 shows results where a fixed 7B supervised model first performs recursive self-critique, followed by evaluation from critic models of varying sizes at each stage. The critics’ judgments are used to train reward models specific to each critic size, which then guide Best-of-N sampling on the same 7B supervised model. The PR metric compares this Best-of-N performance against the oracle Pass@N of this fixed 7B model across different critic sizes.
Recursive self-critiquing benefits weak-to-strong supervision.
Figure 3 demonstrates that when supervised models are larger than the 7B critic model, higher-order critiques generally yield improved performance compared to lower-order critiques. Similarly, Figure 3 shows that when critic models are smaller than the 7B supervised model, higher-order recursive critiques can provide better supervision effectiveness. Both findings support recursive self-critiquing as a promising approach to scalable oversight, particularly in scenarios where humans (as the "weaker model") oversee increasingly capable AI systems (the stronger model).
Direct supervision exhibits superior performance in strong-to-weak settings.
Conversely, both 3 and 3 show that when critic models are stronger than the supervised model, direct critique produces better results than allowing the supervised model to engage in higher-order self-critique first. This asymmetry indicates that self-critique from weaker models may not be effective and can even mislead stronger critics.
6 Discussion and Related Work
@@
RLHF has emerged as the dominant approach in AI alignment, building upon the principle that "verification is easier than generation" (Irving et al., 2018b). However, it utilizes static reward models as proxies for human preferences and thus introduces the risk of reward hacking (Gao et al., 2022; Karwowski et al., 2023). While approaches such as iterative annotation and tool augmentation (Li et al., 2024a; Gou et al., 2024) provide intermediate solutions, they face limitations in supervision capability.
Related Approaches to Scalable Oversight.
As AI capabilities exceed human expertise, humans may no longer provide effective supervision (Amodei et al., 2016). One approach to this challenge is the debate protocol (Irving et al., 2018a), where agents argue for opposing answers (Khan et al., 2024; Michael et al., 2023), though with demonstrated limitations (Kenton et al., 2024a). Unlike debate’s adversarial framework, our approach assumes higher-order critic tasks are inherently easier. Task decomposition methods (Wu et al., 2021) break complex oversight into manageable sub-problems through breadth-first strategies, while our method employs depth-first recursive refinement. Our majority vote baseline builds on self-consistency methods (Wang et al., 2023; Fluri et al., 2023).
@@
The effectiveness of recursive self-critiquing stems from several mechanisms: higher-order criticism shifts attention from details to abstract principles, each critique level provides structured context, and the recursive structure transforms absolute tasks into pairwise judgments, leveraging humans’ advantage in relative assessment (Jones and Inglis, 2015; Kelly et al., 2022). However, our AI experiments show that direct supervision can be superior in strong-to-weak settings, which may be due to the limited critique capabilities of current models (Xi et al., 2024). Future work may focus on enhancing model critique capabilities (Wang et al., 2024a; Yu et al., 2025; Ankner et al., 2024).
@@ 7 Conclusion
This work investigates how to obtain reliable supervision signals when AI capabilities surpass human abilities. Through Human-Human, Human-AI, and AI-AI experiments, we examine the hypothesis that critique of critique is easier than critique and demonstrate that this difficulty relation holds recursively. The experiments suggest a promising pathway for scalable oversight through recursive self-critiquing when direct human evaluation becomes infeasible.
Impact Statement
Our recursive self-critiquing framework offers potential for maintaining effective AI oversight as capabilities surpass human abilities. However, this approach carries risks, including false confidence in oversight effectiveness, vulnerability to adversarial examples. Our experiments also reveal current limitations in AI models’ recursive self-critiquing capabilities, highlighting the need for continued development of models’ self-critique abilities to enhance oversight robustness. We acknowledge these potential impacts and encourage continued research to strengthen scalable oversight methods.
@@ Appendix A Human Experiments Guidelines
This section details the guidelines and quality assurance involved in the Human-Human and Human-AI experiments. We establish consistent and comprehensive guidelines for annotation tasks at different stages across various tasks. Our guidelines emphasize the quality of the reasoning process over accuracy rates, requiring annotators to articulate their thinking process clearly without accessing external references. While accuracy is encouraged, the primary focus is on providing clear, well-reasoned justifications for their decisions. Annotators are instructed to invest their time primarily in analytical thinking, expressing their reasoning in clear, concise, and logically coherent natural language. The guidelines provide suggested formats but maintain flexibility, prioritizing the clear documentation of thought processes over rigid adherence to specific forms [^1] Fixed templates were initially tested but abandoned as annotators reported them to be unflexible and burdensome.. We provide detailed instructions at each stage in the following sections.
@@ A.3 Higher-Order Critique Stage
Examples of high-quality and low-quality higher-order critiques are presented in Tables 9 and 10.